# An impossibility theorem concerning positive involvement in voting

### v4

@@ 3 Impossibility
Smith set 960 2.375 4
Uncovered set 960 2 3
Copeland 960 1.625 3
@@
$\Rightarrow_{d}$
$\mathbb{M}_{2}$
$\Leftarrow_{b}$
$\mathbb{M}_{3}$
@@
$\Rightarrow_{a}$
$\mathbb{M}_{4}$
$\Leftarrow_{d}$
$\mathbb{M}_{5}$
@@
$8$ $8$ $14$ $6$ $7$ $2$ $a$ $a$ $b$ $c$ $c$ $d$ $c$ $d$ $d$ $b$ $d$ $c$ $b$ $c$ $a$ $a$ $b$ $b$ $d$ $b$ $c$ $d$ $a$ $a$ $\mathbf{P}_{1}$ $c$ beats $b$ by $31-14=17$ $a$ beats $c$ by $30-15=15$ $b$ beats $a$ by $29-16=13$ $b$ beats $d$ by $28-17=11$ $d$ beats $c$ by $24-21=3$ $d$ beats $a$ by $23-22=1$
$3$
$d$
$b$
$a$
$c$
$\mathbf{P}_{1\to 2}$
$8$ $8$ $14$ $6$ $7$ $3$ $2$
$a$ $a$ $b$ $c$ $c$ $d$ $d$
$c$ $d$ $d$ $b$ $d$ $b$ $c$
$b$ $c$ $a$ $a$ $b$ $a$ $b$
$d$ $b$ $c$ $d$ $a$ $c$ $a$
$\mathbf{P}_{2}$
$a$ beats $c$ by $33-15=18$
$b$ beats $a$ by $32-16=16$
$c$ beats $b$ by $31-17=14$
$b$ beats $d$ by $28-20=8$
$d$ beats $c$ by $27-21=6$
$d$ beats $a$ by $26-22=4$
@@
$10$ $16$ $1$ $5$ $7$ $16$ $a$ $b$ $b$ $b$ $d$ $d$ $c$ $a$ $a$ $d$ $a$ $c$ $b$ $c$ $d$ $a$ $c$ $b$ $d$ $d$ $c$ $c$ $b$ $a$ $\mathbf{Q}_{2}$ $a$ beats $c$ by $39-16=23$ $b$ beats $a$ by $38-17=21$ $c$ beats $b$ by $33-22=11$ $b$ beats $d$ by $32-23=9$ $d$ beats $c$ by $29-26=3$ $d$ beats $a$ by $28-27=1$
$4$
$b$
$d$
$a$
$c$
$\mathbf{Q}_{2\leftarrow 3}$
$10$ $16$ $1$ $1$ $7$ $16$
@@
$\mathbf{Q}_{3}$
$a$ beats $c$ by $35-16=19$
$b$ beats $a$ by $34-17=17$
$c$ beats $b$ by $33-18=15$
$b$ beats $d$ by $28-23=5$
$a$ beats $d$ by $27-24=3$
$c$ beats $d$ by $26-25=1$
@@
$14$ $4$ $8$ $11$ $2$ $a$ $b$ $b$ $c$ $c$ $d$ $a$ $d$ $b$ $d$ $c$ $c$ $a$ $d$ $b$ $b$ $d$ $c$ $a$ $a$ $\mathbf{R}_{1}$ $c$ beats $b$ by $27-12=15$ $a$ beats $c$ by $26-13=13$ $b$ beats $a$ by $25-14=11$ $b$ beats $d$ by $23-16=7$ $d$ beats $c$ by $22-17=5$ $d$ beats $a$ by $21-18=3$
$2$
$a$
$d$
$c$
$b$
$\mathbf{R}_{1\to 4}$
$16$ $4$ $8$ $11$ $2$
@@
$\mathbf{R}_{4}$
$c$ beats $b$ by $29-12=17$
$a$ beats $c$ by $28-13=15$
$b$ beats $a$ by $25-16=9$
$d$ beats $c$ by $24-17=7$
$b$ beats $d$ by $23-18=5$
$d$ beats $a$ by $21-20=1$
@@
$16$ $9$ $5$ $4$ $3$ $3$ $14$ $a$ $b$ $b$ $c$ $d$ $d$ $d$ $c$ $a$ $d$ $d$ $a$ $b$ $c$ $b$ $d$ $a$ $b$ $c$ $a$ $b$ $d$ $c$ $c$ $a$ $b$ $c$ $a$ $\mathbf{S}_{4}$ $c$ beats $b$ by $37-17=20$ $a$ beats $c$ by $36-18=18$ $b$ beats $a$ by $35-19=16$ $d$ beats $c$ by $34-20=14$ $b$ beats $d$ by $30-24=6$ $d$ beats $a$ by $29-25=4$
$3$
$d$
$a$
$c$
$b$
$\mathbf{S}_{4\leftarrow 5}$
$16$ $9$ $5$ $4$ $3$ $14$
$a$ $b$ $b$ $c$ $d$ $d$
$c$ $a$ $d$ $d$ $b$ $c$
$b$ $d$ $a$ $b$ $a$ $b$
$d$ $c$ $c$ $a$ $c$ $a$
$\mathbf{S}_{5}$
$b$ beats $a$ by $35-16=19$

### v5

@@ 3 Impossibility
Smith set 960 2.375111 4
Uncovered set 960 2.000000 3
Copeland 960 1.625111 3
@@
$\Rightarrow_{d}$ $a$$c$$b$$d$666655551111222233334444 $\mathbb{M}_{2}$ $\Leftarrow_{b}$ $a$$c$$b$$d$666655552222111133334444 $\mathbb{M}_{3}$
@@
$\Rightarrow_{a}$ $a$$c$$b$$d$555544441111333322226666 $\mathbb{M}_{4}$ $\Leftarrow_{d}$ $a$$c$$b$$d$444466661111333322225555 $\mathbb{M}_{5}$
@@
$8$ $8$ $14$ $6$ $7$ $2$
$a$ $a$ $b$ $c$ $c$ $d$
$c$ $d$ $d$ $b$ $d$ $c$
$b$ $c$ $a$ $a$ $b$ $b$
$d$ $b$ $c$ $d$ $a$ $a$
$\mathbf{P}_{1}$
$c$ beats $b$ by $31-14=17$
$a$ beats $c$ by $30-15=15$
$b$ beats $a$ by $29-16=13$
$b$ beats $d$ by $28-17=11$
$d$ beats $c$ by $24-21=3$
$d$ beats $a$ by $23-22=1$
$+$ $3$ $d$ $b$ $a$ $c$ $\mathbf{P}_{1\to 2}$ $=$ $8$ $8$ $14$ $6$ $7$ $3$ $2$ $a$ $a$ $b$ $c$ $c$ $d$ $d$ $c$ $d$ $d$ $b$ $d$ $b$ $c$ $b$ $c$ $a$ $a$ $b$ $a$ $b$ $d$ $b$ $c$ $d$ $a$ $c$ $a$ $\mathbf{P}_{2}$ $a$ beats $c$ by $33-15=18$ $b$ beats $a$ by $32-16=16$ $c$ beats $b$ by $31-17=14$ $b$ beats $d$ by $28-20=8$ $d$ beats $c$ by $27-21=6$ $d$ beats $a$ by $26-22=4$
@@
$10$ $16$ $1$ $5$ $7$ $16$
@@
$\mathbf{Q}_{2}$
$a$ beats $c$ by $39-16=23$
$b$ beats $a$ by $38-17=21$
$c$ beats $b$ by $33-22=11$
$b$ beats $d$ by $32-23=9$
$d$ beats $c$ by $29-26=3$
$d$ beats $a$ by $28-27=1$
$=$ $4$ $b$ $d$ $a$ $c$ $\mathbf{Q}_{2\leftarrow 3}$ $+$ $10$ $16$ $1$ $1$ $7$ $16$ $a$ $b$ $b$ $b$ $d$ $d$ $c$ $a$ $a$ $d$ $a$ $c$ $b$ $c$ $d$ $a$ $c$ $b$ $d$ $d$ $c$ $c$ $b$ $a$ $\mathbf{Q}_{3}$ $a$ beats $c$ by $35-16=19$ $b$ beats $a$ by $34-17=17$ $c$ beats $b$ by $33-18=15$ $b$ beats $d$ by $28-23=5$ $a$ beats $d$ by $27-24=3$ $c$ beats $d$ by $26-25=1$
@@
$14$ $4$ $8$ $11$ $2$
@@
$\mathbf{R}_{1}$
$c$ beats $b$ by $27-12=15$
$a$ beats $c$ by $26-13=13$
$b$ beats $a$ by $25-14=11$
$b$ beats $d$ by $23-16=7$
$d$ beats $c$ by $22-17=5$
$d$ beats $a$ by $21-18=3$
$+$ $2$ $a$ $d$ $c$ $b$ $\mathbf{R}_{1\to 4}$ $=$ $16$ $4$ $8$ $11$ $2$ $a$ $b$ $b$ $c$ $c$ $d$ $a$ $d$ $b$ $d$ $c$ $c$ $a$ $d$ $b$ $b$ $d$ $c$ $a$ $a$ $\mathbf{R}_{4}$ $c$ beats $b$ by $29-12=17$ $a$ beats $c$ by $28-13=15$ $b$ beats $a$ by $25-16=9$ $d$ beats $c$ by $24-17=7$ $b$ beats $d$ by $23-18=5$ $d$ beats $a$ by $21-20=1$
@@
$16$ $9$ $5$ $4$ $3$ $3$ $14$
$a$ $b$ $b$ $c$ $d$ $d$ $d$
$c$ $a$ $d$ $d$ $a$ $b$ $c$
$b$ $d$ $a$ $b$ $c$ $a$ $b$
$d$ $c$ $c$ $a$ $b$ $c$ $a$
$\mathbf{S}_{4}$
$c$ beats $b$ by $37-17=20$
$a$ beats $c$ by $36-18=18$
$b$ beats $a$ by $35-19=16$
$d$ beats $c$ by $34-20=14$
$b$ beats $d$ by $30-24=6$
$d$ beats $a$ by $29-25=4$
$=$ $3$ $d$ $a$ $c$ $b$ $\mathbf{S}_{4\leftarrow 5}$ $+$ $16$ $9$ $5$ $4$ $3$ $14$ $a$ $b$ $b$ $c$ $d$ $d$ $c$ $a$ $d$ $d$ $b$ $c$ $b$ $d$ $a$ $b$ $a$ $b$ $d$ $c$ $c$ $a$ $c$ $a$ $\mathbf{S}_{5}$ $b$ beats $a$ by $35-16=19$